A cyclist traveled 9 kilometers per hour faster than an in-line skater. In the time it took the cyclist to travel 45 kilometers, the skater had gone 30 kilometers. Find the speed of the skater.
Answer by jorel555(1290)

Let x=speed of skater, in kilometers per hour;
x+9=speed of the cyclist; thus:
45/(x+9)=30/(x)
45x=30x+270
15x=270
x=18
x+9=27.
Let's check:
45/27=1.6666666
30/18=1.6666666 !!!!!!!!